Could you please help me out with this problem?
Solve using addition and multiplication together...
3y-8= -1+2y
Thanks!!!!!

Found 2 solutions by checkley77, rfer:
Answer by checkley77(12844)

3y-8=-1+2y
3y-2y=-1+8
y=7 ans.
Proof:
3*7-8=-1+2*7
21-8=-1+14
13=13

Answer by rfer(16322)

3y-8=-1+2y
add 8 to both sides
3y-8+8=-1+8+2y
3y=+7+2y
subtract 2y from both sides
3y-2y=7+2y-2y
y=7